OpenStudy (anonymous):

One number is 9 less than three times a second number. If the first number plus twice the second number is 16, find the two number?

OpenStudy (anonymous):

Say x = number 1 y = number 2 x = 3y - 9 x + 2y = 16 You're trying to find y, so, change the equations a bit. x = 16 - 2y Then you can plug this x value into the other equation, and you get 16 - 2y = 3y - 9 Combine these 5y = 25 y = 5 That make sense?

OpenStudy (anonymous):

Nicely illustrated. Thanks a lot champ